In a certain large state university, older students get to choose their classes before younger students; that often means that freshmen don't get to take the classes they want. Jacob wants to take a really popular course in Russian literature, but his adviser says there's only a 30% chance that it will be available when he gets to choose his classes. Jacob wonders if his chances are actually better than 30% since in the past few years, 73 of 200 freshmen who requested the course were able to get in. Do these data provide good evidence that freshmen have a better than 30% chance of getting the course if they ask for it?

(a)

What is the null hypothesis?

correctH0: p = 0.30H0: p < 0.30    H0: p > 0.30H0: p ≠ 0.30Incorrect: Your answer is incorrect.

(b)

What is the alternative hypothesis?

Ha: p = 0.30Ha: p < 0.30    Ha: p > 0.30Ha: p ≠ 0.30Correct: Your answer is correct.

(c)

What is the value of ?

=

(d)

What is the value of z? (Round your answer to two decimal places.)

z =

(e)

What is the P-value for this test? (Use a table or technology. Round your answer to four decimal places.)
Answer #1

Solution :

This is the right tailed test .

The null and alternative hypothesis is

H0 : p = 0.30

Ha : p > 0.30

\hat p = x / n = 73 / 200 = 0.365

Test statistic = z

= \hat p - P0 / [\sqrtP0 * (1 - P0 ) / n]

= 0.365 - 0.30 / [\sqrt(0.30 * 0.70) / 200]

z = 2.01

P-value = 0.0224